Susie started with 100g of calcium carbonate. At the end of the reaction there was 56g left in the test tube. Calculate the mass of carbon dioxide made


1
Expert's answer
2021-03-22T06:27:37-0400

100-56 = 44g

1mol CaCO3 = 1molCO2

Molar Mass CaCO3 = 100g/mol

44g CaCO3 = 44g/100g/mol = 0.44mol

0.44mol CO2 made

Molar Mass of CO2 = 44g/mol

Mass of 0.44mol CO2 = 0.44×44

= 19.36g CO2 made